# Addressing Polarization and Unfairness in Performative Prediction

### v1

@@ Appendix D Proofs for the Theoretical Results
D.1 Proof of Proposition 5.1
@@
(First order optimality condition of convex functions [41]).
@@
$=$ $~{}-\alpha(1-\alpha)[\ell(\boldsymbol{\theta}_{1})-\ell^{2}(\boldsymbol{\theta}_{1})]^{2}+(1-\alpha)\ell^{2}(\boldsymbol{\theta}_{1})+\alpha\ell^{2}(\boldsymbol{\theta}_{2})$
$\leq$ $~{}(1-\alpha)\ell^{2}(\boldsymbol{\theta}_{1})+\alpha\ell^{2}(\boldsymbol{\theta}_{2}),$
@@
Under conditions in Lemma 2.4, the loss aware objective $\mathcal{L}_{\textsf{fair}}$ in Eqn (7) and (8) are $\tilde{\beta}$-jointly smooth, where $\tilde{\beta}=(2\rho\overline{\ell}+1)\beta$.
@@
$~{}\|\nabla_{\boldsymbol{\theta}}g(\boldsymbol{\theta};Z)-\nabla_{\boldsymbol{\theta}}g(\boldsymbol{\theta}^{\prime};Z)\|_{2}\leq 2\beta\max\{\ell(\boldsymbol{\theta};Z),\ell(\boldsymbol{\theta}^{\prime};Z)\}\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{2},$ (17)
$~{}\|\nabla_{\boldsymbol{\theta}}g(\boldsymbol{\theta};Z)-\nabla_{\boldsymbol{\theta}}g(\boldsymbol{\theta};Z^{\prime})\|_{2}\leq 2\beta\max\{\ell(\boldsymbol{\theta};Z),\ell(\boldsymbol{\theta};Z^{\prime})\}\|Z-Z^{\prime}\|_{2},$
which implies $(2\rho\overline{\ell}+1)\beta$-joint smoothness of $\mathcal{L}_{\textsf{fair}}$ in Eqn (7) and (8). ∎
@@
$\boldsymbol{\theta}^{(t+1)}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}=\arg\min_{\boldsymbol{\theta}}\sum_{s}p_{s}^{(t)}(1+\rho\mathcal{L}_{s}(\boldsymbol{\theta}^{(t)});\mathcal{D}^{(t-1)}_{s})\mathcal{L}_{s}(\boldsymbol{\theta});\mathcal{D}^{(t)}_{s},$ (18)
@@
At a high level, $(1+\rho\mathcal{L}_{s}(\boldsymbol{\theta}^{(t)});\mathcal{D}^{(t-1)}_{s})$ acts like a scalar that may scale the strong convexity and joint smoothness coefficient.
Since $\mathcal{L}_{s}(\boldsymbol{\theta}^{(t)});\mathcal{D}^{(t-1)}_{s})\geq 0$, we know the $\gamma$-strong convexity part holds.
Similarly, since $\mathcal{L}_{s}(\boldsymbol{\theta}^{(t)});\mathcal{D}^{(t-1)}_{s})\leq\overline{\ell}$, we know the $(1+\rho\overline{\ell})$-joint smoothness holds. ∎
@@
Based on the above lemmas, we present a result that bounds the fair-aware loss minimization solution distance from two different input distributions, where [4] and [5] derived similar bounds in the PP cases without fairness penalties.
@@
$\boldsymbol{\theta}_{\textsf{fair}}^{*}=G(\mathcal{D};\mathcal{L}_{\textsf{fair}}):=\arg\min_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}),~{}~{}~{}\tilde{\boldsymbol{\theta}}_{\textsf{fair}}^{*}=G(\tilde{\mathcal{D}};\mathcal{L}_{\textsf{fair}}):=\arg\min_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\tilde{\mathcal{D}}),$ (19)
@@
$\|\boldsymbol{\theta}_{\textsf{fair}}^{*}-\tilde{\boldsymbol{\theta}}_{\textsf{fair}}^{*}\|_{2}\leq\frac{\tilde{\beta}}{\gamma}\mathcal{W}_{1}(\mathcal{D},\tilde{\mathcal{D}}).$ (20)
@@
$-\gamma\|G(\mathcal{D};\mathcal{L}_{\textsf{fair}})-G(\tilde{\mathcal{D}};\mathcal{L}_{\textsf{fair}})\|_{2}^{2}\geq(G(\mathcal{D};\mathcal{L}_{\textsf{fair}})-G(\tilde{\mathcal{D}};\mathcal{L}_{\textsf{fair}}))^{\top}\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}).$ (21)
@@
$(G(\mathcal{D};\mathcal{L}_{\textsf{fair}})-G(\tilde{\mathcal{D}};\mathcal{L}_{\textsf{fair}}))^{\top}\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D})\geq-(2\rho\overline{\ell}+1)\beta\mathcal{W}_{1}(\mathcal{D},\tilde{\mathcal{D}}),$ (24)
@@
$-\gamma\|G(\mathcal{D};\mathcal{L}_{\textsf{fair}})-G(\tilde{\mathcal{D}};\mathcal{L}_{\textsf{fair}})\|_{2}^{2}\geq-(2\rho\overline{\ell}+1)\beta\mathcal{W}_{1}(\mathcal{D},\tilde{\mathcal{D}}),$ (25)
@@
(Bounding the fixed point distribution distance using parameter distance [5])
@@
Then using the results in Lemma D.2, D.3, D.5, D.4, D.6, we can see that
@@
([5]) Under conditions in Lemma 2.4, given a decision model with parameter $\boldsymbol{\theta}$, denote the parameter returned by Delayed Deployment Scheme as $\hat{\mathcal{D}}_{\theta}:=\text{T}^{DL}(\boldsymbol{\theta};\mathcal{D})$ and the fixed point distribution of $\boldsymbol{\theta}$ as $\mathcal{D}_{\boldsymbol{\theta}}$, then
@@
 [^1] Due to different notations, we present the proof in [5] to help the readers.
@@
The proof of this part referenced [5] Theorem 4. ∎
@@
The proof of this part referenced [5] Theorem 8. ∎
D.3 Proof of Theorem B.3
@@
[…]
For given a convergence radius $\delta\in(0,1)$, take $n_{t}=O\left(\frac{\log(t/p)}{(\epsilon(1+\tilde{\beta}/\gamma)\delta)^{m}}\right)$ samples at $t$. If $12\epsilon(1+\tilde{\beta}/\gamma)<$ 1, then with probability $1-p$, the iterates of fair-RERM are within a radius $\delta$ of the Fair-PS pair for $t\geq(1-2\epsilon(1+\tilde{\beta}/\gamma)O(\log(1/\delta))$.
@@
[…]
@@
i.e., with probability at least $1-\sum_{t=1}^{\infty}1-p$ we have that for each time step $t$, it holds that
@@
The proof of this part referenced Theorem 5 in [5]
D.4 Proof of the Fairness Guarantee
Discussion of the intricate nature of proving the effectiveness of $\rho$ under the general PP setting. We first note that it is non-trivial to find out whether fairness at PS solution increases as $\rho$ increases under the general setting where both the group fractions and the group-wise feature-label distributions change. Consider an example where the agents in the minority group are influenced by the majority group, i.e., when their group fraction becomes less than a threshold, their tastes are shaped by the mainstream culture. Then they can change their feature-label distributions to be more similar to the one of the majority group. This will possibly make the loss disparity smaller even when $\rho$ becomes larger. However, in majority of previous work on the retention of recommendation system [2, 12], the group-wise distributions are assumed to be static.
@@ D.4 Proof of the Fairness Guarantee
[…]
@@
    Otherwise Eqn. (51) would be smaller than 00. We have by assumption $L_{a,\rho_{2}}^{t}>L_{a,\rho_{1}}^{t}>L_{b,\rho_{1}}^{t}>L_{b,\rho_{2}}^{t}$. Then $L_{a,\rho_{1}}^{t}+L_{a,\rho_{2}}^{t}>L_{b,\rho_{1}}^{t}+L_{b,\rho_{2}}^{t}$. Next, noticing that $L_{b,\rho_{2}}^{t}-L_{b,\rho_{1}}^{t}<0$ and $L_{a,\rho_{2}}^{t}-L_{a,\rho_{1}}^{t}>0$, then if we divide $L_{a,\rho_{1}}^{t}+L_{a,\rho_{2}}^{t}$ at both sides of Eqn. (52), then we will get:
@@
D.5 Discussion on Performative Optimal (PO) Solutions

### v2

@@ C.2 Gaussian mean estimation with multiple groups
C.5 Gaussian mean estimation Using group loss variance
In Example 4.2, we already prove that group loss variance as a fair penalty term violates the strong convexity assumption. In this section, we show its unconvergence while performing the Gaussian mean estimation task under the same setting mentioned in Section 6 when $\rho=\{0.5,1.5,2.0\}$. The dynamics of $\|\theta_{t}-\theta_{t-1}\|$ as well as performative loss are shown as follows. Fig. 7 shows the failure of convergence, which is contradicted to Fig. 9(a) where we show the convergence of Fair-RERM.
C.6 K-delayed RERM on Credit Data
We examine the effectiveness of k-delayed RERM schema on Credit Data (creditdata) which is the same dataset used by brown2020performative. Specifically, we use $k=3$ and $\epsilon=1$ to form equal-sized agent groups uniformly in the whole population to let each subgroup best respond with different speed (Example 1 in brown2020performative). We also let the retention dynamic be k-delayed, adding additional challenges compared to the previous work. We examine $\rho\in\{0.3,0.7\}$ while all other settings are the same as the Credit Data experiment in Section 6.
C.7 Complementary results on Equal Opportunity and Demographic Parity
Though group-wise loss disparity has been a widely used fairness metric for both regression and classification tasks, we also provide a complementary experimental result on how our fairness mechanism influences Equal opportunity (hardt2016equality). Specifically, we visualize the evolution of the true positive rate disparity between two groups until convergence to the Fair-PS point for the two classification tasks, and we visualize the evolution of the positive rate disparity. The results are in Fig. 8, which empirically demonstrate the effectiveness of all fairness mechanisms in improving equal opportunity and demographic parity.
@@ Appendix D Proofs for the theoretical results
D.1 Proof of Lemma 2.4
Proof.
The first part of Lemma 2.4 is just Thm. 8 in brown2020performative. For the second part, we prove that when (ii) and (iii) are satisfied, there exists a non-convex loss function $\ell$ where the iterative algorithms fail to converge by constructing an example:
Consider the cubic loss $\ell((x,y);\theta)=\beta y\theta^{3}$, for $\theta\in[-1,1]$. This objective is $\beta$-jointly smooth but not convex. Let the distribution of $Y$ according to $\text{T}(\boldsymbol{\theta},\mathcal{D})$ be a point mass at $\epsilon\theta$, and let the distribution of $X$ be invariant. Then this distribution is $\epsilon$-sensitive.
Then the decoupled performative risk has the following form: $\text{DPR}(\theta,\varphi)=\epsilon\beta\theta^{3}\varphi$. Then if we initialize RRM at any point other than 0 with any starting distribution and $\boldsymbol{\theta}^{(0)}>0$, the procedure generates the sequence of iterates $\dots,-1,1,-1,1,\dots$, thus failing to converge. ∎
D.2 Propositions 3.1 and 3.2
We can prove Prop. 3.1 and Prop. 3.2 by giving two examples where the polarization effect happens or group-wise loss disparity increases.
Example D.1 (Polarization effects of $\boldsymbol{\theta}^{\text{PS}}$).
Consider individuals from two groups $a,b$ whose participation in an ML system depends on their perceived loss, i.e., group fraction $p_{s}^{(t)}$ changes based on the deployed ML model. Suppose individuals from group $a$ and $b$ have fixed data $Z=z_{a}=1$ and $Z=z_{b}=0$, respectively. Their initial group fractions are $p_{a}^{(0)}=0.7,p_{b}^{(0)}=0.3$. Consider conventional RRM with mean squared error $\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t)})=\sum_{s\in\{a,b\}}p_{s}^{(t)}(\boldsymbol{\theta}-z_{s})^{2}$ and transition T satisfies $p_{a}^{(t+1)}=p_{a}^{(t)}+0.1\cdot\left[\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{b}^{(t)})-\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{a}^{(t)})\right],p_{b}^{(t+1)}=1-p_{a}^{(t+1)}$ (i.e., group fraction decreases if the group has higher loss). Next, consider $\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{(t)})=p_{a}^{(t)}(\boldsymbol{\theta}-1)^{2}+p_{b}^{(t)}\boldsymbol{\theta}^{2}$. Minimizing the loss results in $\boldsymbol{\theta}^{(t)}=\frac{p_{a}^{(t)}}{p_{a}^{(t)}+p_{b}^{(t)}}$. This means when $p_{a}^{(t)}>0.5$, $\boldsymbol{\theta}^{(t)}>0.5$ and $\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{b}^{(t)})>\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{a}^{(t)})$, which further results in the increase of $p_{a}^{(t+1)}$. Suppose $\boldsymbol{\theta}^{(0)}=0.7$ is the risk minimizer with respect to the initial distribution. Thus, $p_{a}^{(t)}$ (resp. $p_{b}^{(t)}$) monotonically increases (resp. decreases) in $t$, and the PS solution solely contains people from group $a$ (see Fig. 1).
Example D.2 (Exacerbated group-wise loss disparity).
Consider individuals from two groups $a,b$ who are subject to certain ML decisions; each has an initial feature $X\sim\text{Uniform}(0,1)$ and a binary label $Y=\mathbf{1}(X\geq 0.5)\in\{0,1\}$. Suppose the population has fixed group fractions $p_{a}=0.7,p_{b}=0.3$, but individuals may strategically manipulate their features to increase their chances of receiving favorable ML decisions, i.e., distribution of each group $\mathcal{D}_{s}^{(t)}$ changes based on ML model. Following the individual response model in (perdomo_performative_2021), we suppose individuals from two groups have certain effort budgets $\eta_{a}=0.2,\eta_{b}=0.01$ that can be used to manipulate their features $X_{s}^{(t+1)}=X_{s}+\eta_{s}\cdot\boldsymbol{\theta}^{(t)}$ without changing $Y$. Consider RRM iterative algorithm where the model parameter $\boldsymbol{\theta}^{(t)}$ is updated to maximize prediction accuracy $\Pr(\mathbf{1}(X^{(t)}\geq\boldsymbol{\theta}^{(t)})=Y)$. The initial model $\boldsymbol{\theta}^{(0)}=0.5$ which is the risk minimizer with respect to the initial distribution. Then at $t=1$, all individuals from group $a$ has features increasing $0.1$ but their labels are unchanged. Because group $a$ is the majority group, the risk minimizer at next round $\boldsymbol{\theta}^{(1)}$ increases to $0.6$. $\boldsymbol{\theta}^{(t)}$ will then keep increasing over time until its convergence to PS solution $\boldsymbol{\theta}^{\text{PS}}=0.625$ (see Fig. 1). However, at the PS solution, the accuracy for group $a$ is 1 but is only $0.88$ for group $b$.
D.3 Proof of Proposition 5.1
@@
(First order optimality condition of convex functions (Bubeck2015book)).
@@
$=$ $~-\alpha(1-\alpha)[\ell(\boldsymbol{\theta}_{1})-\ell(\boldsymbol{\theta}_{2})]^{2}+(1-\alpha)\ell^{2}(\boldsymbol{\theta}_{1})+\alpha\ell^{2}(\boldsymbol{\theta}_{2})$
$\leq$ $~(1-\alpha)\ell^{2}(\boldsymbol{\theta}_{1})+\alpha\ell^{2}(\boldsymbol{\theta}_{2})$
@@
Under conditions in Lemma 2.4, the loss aware objective $\mathcal{L}_{\textsf{fair}}$ in Eqn (7) and (8) are $\tilde{\beta}-$jointly smooth, where $\tilde{\beta}=(2\rho\overline{\ell}+1)\beta+2\rho\tilde{\ell}^{2}$.
@@
$~\|\nabla_{\boldsymbol{\theta}}g(\boldsymbol{\theta};Z)-\nabla_{\boldsymbol{\theta}}g(\boldsymbol{\theta}^{\prime};Z)\|_{2}\;\leq 2\beta\max\{\|\ell(\boldsymbol{\theta};Z)\|,\|\ell(\boldsymbol{\theta}^{\prime};Z)\|\}\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{2}\;+2\,\tilde{\ell}^{\,2}\;\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{2},$ (18)
$~\|\nabla_{\boldsymbol{\theta}}g(\boldsymbol{\theta};Z)-\nabla_{\boldsymbol{\theta}}g(\boldsymbol{\theta};Z^{\prime})\|_{2}\;\leq 2\beta\max\{\|\ell(\boldsymbol{\theta};Z)\|,\|\ell(\boldsymbol{\theta};Z^{\prime})\|\}\|Z-Z^{\prime}\|_{2}\;+2\,\tilde{\ell}^{2}\;\|Z-Z^{\prime}\|_{2}.$
which implies $\left((2\rho\overline{\ell}+1)\beta+2\rho\tilde{\ell}^{2}\right)$-joint smoothness of $\mathcal{L}_{\textsf{fair}}$ in Eqn (7) and (8). ∎
@@
$\boldsymbol{\theta}^{(t+1)}=\arg\min_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}=\arg\min_{\boldsymbol{\theta}}\sum_{s}p_{s}^{(t)}(1+\rho\mathcal{L}_{s}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)}_{s})\mathcal{L}_{s}(\boldsymbol{\theta});\mathcal{D}^{(t)}_{s}$, (19)
@@
At a high level, $(1+\rho\mathcal{L}_{s}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)}_{s})$ acts like a scalar that may scale the strong convexity and joint smoothness coefficient.
Since $\mathcal{L}_{s}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)}_{s})\geq 0$, we know the $\gamma$-strong convexity part holds.
Similarly, since $\mathcal{L}_{s}(\boldsymbol{\theta}^{(t)};\mathcal{D}^{(t-1)}_{s})\leq\overline{\ell}$, we know the $(1+\rho\overline{\ell})$-joint smoothness holds. ∎
@@
Based on the above lemmas, we present a result that bounds the fair-aware loss minimization solution distance from two different input distributions, where (perdomo_performative_2021) and (brown2020performative) derived similar bounds in the PP cases without fairness penalties.
@@
[…]
@@
[…]
@@
$\gamma\|G(\mathcal{D};\mathcal{L}_{\textsf{fair}})-G(\tilde{\mathcal{D}};\mathcal{L}_{\textsf{fair}})\|_{2}^{2}\geq(G(\mathcal{D};\mathcal{L}_{\textsf{fair}})-G(\tilde{\mathcal{D}};\mathcal{L}_{\textsf{fair}}))^{\top}\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\textsf{fair}}(\boldsymbol{\theta};\mathcal{D}).$ (22)
@@
[…]
@@
[…]
@@
(Bounding the fixed point distribution distance using parameter distance (brown2020performative))
@@
Then using the results in Lemma D.4, D.5, D.7, D.6, D.8, we can see that
@@
((brown2020performative)) Under conditions in Lemma 2.4, given a decision model with parameter $\boldsymbol{\theta}$, denote the parameter returned by Delayed Deployment Scheme as $\hat{\mathcal{D}}_{\theta}:=\text{T}^{DL}(\boldsymbol{\theta};\mathcal{D})$ and the fixed point distribution of $\boldsymbol{\theta}$ as $\mathcal{D}_{\boldsymbol{\theta}}$, then
@@
 [^2] Due to different notations, we present the proof in (brown2020performative) to help the readers.
@@
The proof of this part referenced (brown2020performative) Theorem 4. ∎
@@
The proof of this part referenced (brown2020performative) Theorem 8. ∎
D.5 Proof of Theorem B.3
@@
Theorem B.3 (Fair RERM Convergence) Under conditions in Lemma 2.4, suppose $\exists\alpha>1,\mu>0$ such that $\int_{\mathbb{R}^{m}}e^{\mu|x|^{\alpha}}Zdx$ is finite $\forall Z\in\triangle(\mathcal{Z})$.
For given a convergence radius $\delta\in(0,1)$, take $n_{t}=O\left(\frac{\log(t/p)}{(\epsilon(1+\tilde{\beta}/\gamma)\delta)^{m}}\right)$ samples at $t$. If $2\epsilon(1+\tilde{\beta}/\gamma)<1$, then with probability 1p1-p, the iterates of fair-RERM are within a radius $\delta$ of the Fair-PS pair for $t\geq(1-2\epsilon(1+\tilde{\beta}/\gamma)O(\log(1/\delta))$.
@@
By Theorem 2 of (fournier_hal-00915365), since $\exists\alpha>1,\mu>0$ such that $\int_{\mathbb{R}^{m}}e^{\mu|x|^{\alpha}}Zdx$ is finite $\forall Z\in\triangle(\mathcal{Z})$, then if $n_{t}=\mathcal{O}\left(\frac{1}{\epsilon(1+\tilde{\beta}/\gamma)^{m}}\log(t/p)\right)$, where $m$ is the dimension of the sample, we have $\mathcal{W}_{1}(\mathcal{D},\hat{\mathcal{D}}^{n})\geq\epsilon(1+\tilde{\beta}/\gamma)\delta$ with probability at most $\frac{6p}{\pi^{2}t^{2}}$.
@@
i.e., with probability at least 1p1-p we have that for each time step $t$, it holds that
@@
The proof of this part referenced Theorem 5 in (brown2020performative)
D.6 Proof of the fairness guarantee
Discussion of the intricate nature of proving the effectiveness of $\rho$ under the general PP setting. We first note that it is non-trivial to find out whether fairness at PS solution increases as $\rho$ increases under the general setting where both the group fractions and the group-wise feature-label distributions change. Consider an example where the agents in the minority group are influenced by the majority group, i.e., when their group fraction becomes less than a threshold, their tastes are shaped by the mainstream culture. Then they can change their feature-label distributions to be more similar to the one of the majority group. This will possibly make the loss disparity smaller even when $\rho$ becomes larger. However, in majority of previous work on the retention of recommendation system (Zhang_2019_Retention; duchi2018learning), the group-wise distributions are assumed to be static.
@@ D.6 Proof of the fairness guarantee
    $\Leftrightarrow$~ $p_{a}\cdot\left(L_{a,\rho_{2}}^{t}+L_{a,\rho_{1}}^{t}\right)\cdot\left(L_{a,\rho_{2}}^{t}-L_{a,\rho_{1}}^{t}\right)+p_{b}\cdot\left(L_{b,\rho_{2}}^{t}+L_{b,\rho_{1}}^{t}\right)\cdot\left(L_{b,\rho_{2}}^{t}-L_{b,\rho_{1}}^{t}\right)<0$
@@
    Otherwise Eqn. (52) would be smaller than 0. We have by assumption $L_{a,\rho_{2}}^{t}>L_{a,\rho_{1}}^{t}>L_{b,\rho_{1}}^{t}>L_{b,\rho_{2}}^{t}$. Then $L_{a,\rho_{1}}^{t}+L_{a,\rho_{2}}^{t}>L_{b,\rho_{1}}^{t}+L_{b,\rho_{2}}^{t}$. Next, noticing that $L_{b,\rho_{2}}^{t}-L_{b,\rho_{1}}^{t}<0$ and $L_{a,\rho_{2}}^{t}-L_{a,\rho_{1}}^{t}>0$, then if we divide $L_{a,\rho_{1}}^{t}+L_{a,\rho_{2}}^{t}$ at both sides of Eqn. (53), then we will get:
@@
D.7 Proof of Theorem 5.5
Theorem 5.5 Denote $p_{s}^{\text{PS}}$ as the fraction of group $s$ at $\mathcal{D}^{\text{PS}}$ under retention dynamics. We assume $\mathcal{L}(\boldsymbol{\theta};\mathcal{D})$ is twice continuously differentiable. For sufficiently small $\rho>0$, we know $\Delta_{\mathcal{L}}^{\text{PS}}-\Delta_{\mathcal{L},\text{fair}}^{\text{PS}}(\rho)=2\rho\,p_{a}^{\text{PS}}p_{b}^{\text{PS}}\,\Delta_{\mathcal{L}}^{\text{PS}}\cdot\boldsymbol{v}_{\text{fair}}^{\top}H^{-1}\boldsymbol{v}_{\text{fair}}+\mathcal{O}(\rho^{2})$ where $\boldsymbol{v}_{\text{fair}}=\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}_{a})-\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}_{b}),H=\nabla_{\boldsymbol{\theta}}^{2}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}^{\text{PS}}).$
Proof.
Under fixed $\mathcal{D}_{s}$, the Fair-PS solution $\boldsymbol{\theta}_{\rho}:=\boldsymbol{\theta}^{\text{PS}}_{\text{fair}}(\rho)$ minimizes the fixed objective:
$\mathcal{L}_{\text{fair}}(\boldsymbol{\theta},\rho)=\mathcal{L}(\boldsymbol{\theta};\mathcal{D}^{\text{PS}})+\rho\sum_{s\in\{a,b\}}p_{s}^{\text{PS}}\left[\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s})\right]^{2}.$
Define $F(\boldsymbol{\theta},\rho):=\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{fair}}(\boldsymbol{\theta},\rho)$, so $F(\boldsymbol{\theta}_{\rho},\rho)=0$ and $\boldsymbol{\theta}_{0}=\boldsymbol{\theta}_{\text{PS}}$.
By the implicit function theorem:
$\frac{d\boldsymbol{\theta}_{\rho}}{d\rho}=-\left(\nabla_{\boldsymbol{\theta}}^{2}\mathcal{L}_{\text{fair}}(\boldsymbol{\theta}_{\rho},\rho)\right)^{-1}\cdot\nabla_{\boldsymbol{\theta}}P(\boldsymbol{\theta}_{\rho}),$
where the penalty term is:
$P(\boldsymbol{\theta}):=\sum_{s\in\{a,b\}}p_{s}^{\text{PS}}\left[\mathcal{L}(\boldsymbol{\theta};\mathcal{D}_{s})\right]^{2}.$
At $\rho=0$, this gives:
$\left.\frac{d\boldsymbol{\theta}_{\rho}}{d\rho}\right|_{\rho=0}=-H^{-1}\nabla_{\boldsymbol{\theta}}P(\boldsymbol{\theta}^{\text{PS}}).$
Now compute the gradient of the penalty:
$\nabla_{\boldsymbol{\theta}}P(\boldsymbol{\theta}^{\text{PS}})=2\sum_{s\in\{a,b\}}p_{s}^{\text{PS}}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}_{s})\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}_{s}).$
Using the first-order condition for $\boldsymbol{\theta}^{\text{PS}}$,
$p_{a}^{\text{PS}}\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}_{a})+p_{b}^{\text{PS}}\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}_{b})=0,$
which implies:
$\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}_{b})=-\frac{p_{a}^{\text{PS}}}{p_{b}^{\text{PS}}}\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta}^{\text{PS}};\mathcal{D}_{a}).$
Substituting yields:
$\nabla_{\boldsymbol{\theta}}P(\boldsymbol{\theta}^{\text{PS}})=2p_{a}^{\text{PS}}p_{b}^{\text{PS}}\Delta_{\mathcal{L}}^{\text{PS}}\cdot\boldsymbol{v}_{\text{fair}}.$
So,
$\left.\frac{d\boldsymbol{\theta}_{\rho}}{d\rho}\right|_{\rho=0}=-2p_{a}^{\text{PS}}p_{b}^{\text{PS}}\Delta_{\mathcal{L}}^{\text{PS}}\cdot H^{-1}\boldsymbol{v}_{\text{fair}}.$
Next, we compute the derivative of the loss disparity:
$\left.\frac{d}{d\rho}\Delta\mathcal{L}(\boldsymbol{\theta}_{\rho})\right|_{\rho=0}=\boldsymbol{v}_{\text{fair}}^{\top}\left.\frac{d\boldsymbol{\theta}_{\rho}}{d\rho}\right|_{\rho=0}=-2p_{a}^{\text{PS}}p_{b}^{\text{PS}}\Delta_{\mathcal{L}}^{\text{PS}}\cdot\boldsymbol{v}_{\text{fair}}^{\top}H^{-1}\boldsymbol{v}_{\text{fair}}.$
Finally, apply a first-order Taylor expansion:
$\Delta_{\mathcal{L},\text{fair}}^{\text{PS}}(\rho)=\Delta_{\mathcal{L}}^{\text{PS}}+\rho\left.\frac{d}{d\rho}\Delta\mathcal{L}(\boldsymbol{\theta}_{\rho})\right|_{\rho=0}+\mathcal{O}(\rho^{2}),$
$\Delta_{\mathcal{L}}^{\text{PS}}-2\rho\,p_{a}^{\text{PS}}p_{b}^{\text{PS}}\Delta_{\mathcal{L}}^{\text{PS}}\cdot\boldsymbol{v}_{\text{fair}}^{\top}H^{-1}\boldsymbol{v}_{\text{fair}}+\mathcal{O}(\rho^{2}),$
which completes the proof. ∎
D.8 Discussion on Performative Optimal (PO) Solutions